# Finite temperature detection of quantum critical points via internal quantum teleportation

### v1

@@ IV.2 The $\gamma$ transition
In Figs. 18 and 19 we plot, respectively, $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ as functions of $\gamma$ while fixing $\lambda=1.5$.
@@ V Discussion
The standard approach to study quantum phase transitions also makes use of those correlation functions or of quantities that depend on them (the magnetic susceptibility, for instance). The QCPs are detected at $T=0$ by discontinuities in the $n$-th order derivatives of those quantities at the QCPs. Quantum information theory quantities, such as bipartite or multipartite entanglement and quantum discord, are also useful to spotlight QCPs wu04 ; oli06 ; dil08 ; sar08 ; fan14 ; gir14 ; wer10 ; wer10b . These quantities are also functions of the correlation functions of the system, the same ones we need to theoretically apply the present and the method of Ref. pav23 .
As already pointed out in Ref. pav23 , some of the standard quantities used to detect QCPs (magnetization or magnetic susceptibility, for instance) become less effective to correctly identify the QCPs with finite $T$ data wer10b . And some quantum information theory quantities, in particular the entanglement of formation woo98 , are zero at and around the QCPs as we increase $T$. These quantities cannot be employed to spotlight QCPs after a certain temperature threshold wer10b . On the other hand, a specific measure of non-classical correlations between two systems, namely, quantum discord oll01 ; hen01 , is extremely robust to spotlight QCPs at finite $T$ wer10b . The present internal teleportation based tools to detect QCPs as well as the ones of Ref. pav23 share the quantum discord robustness to detect QCPs with finite $T$ data. In addition to that, and contrary to quantum discord, the present tools and the ones of Ref. pav23 do not become computationally intractable if we deal with high spin systems (high dimensional Hilbert spaces) pav23 .
@@
Furthermore, to execute the teleportation protocol, and thus the present proposal, we need to be able to implement Bell state measurements and local unitary operations on the qubits that belong to a spin chain. A Bell state measurement is a joint measurement involving two qubits that aims to project the two qubits onto one of the four Bell states. This measurement can be divided into the following sequence of three steps nie00 . First, we implement a controlled-not (CNOT) gate on the pair of spins we will be measuring, now called control and target qubits. The CNOT is a two-qubit (non-local) gate whose “truth table” in the standard basis is such that it flips the spin of the target qubit if, and only if, the control qubit is given by $|1\rangle$. The standard basis can be, for instance, the eigenstates of the spin operator $\sigma^{z}$. Second, we implement a Hadamard gate on the control qubit of the CNOT gate. This is a single (local) qubit gate whose truth table is $(|0\rangle+(-)|1\rangle)/\sqrt{2}\rightarrow|0\rangle(|1\rangle)$. Third, we measure the two qubits in the standard (computational) basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Each one of the four previous outcomes in the standard basis implies that we projected the two qubits onto one of the four Bell states nie00 . The Bell measurement is Alice’s job while Bob’s job is to measure his qubit after implementing the appropriate local operation (local gate) as given by Eqs. (8)-(11). All these gates, the ones needed by Alice and Bob, are already being implemented in spin chain-like systems using state of the art techniques ron15 ; bra19 ; xie19 ; noi22 ; xue22 ; mad22 ; xie22 . We believe that in the near future they can be implemented in the sequence described above to execute the teleportation protocol and consequently the present proposal.
@@
For instance, using superconducting qubits to realize a spin chain, it is already possible to experimentally prepare tens of spins simultaneously with coherence times around $100\mu s$. Also, single and two-qubit gates in this platform can be implemented within $10ns$ to $100ns$, leading to the execution of $10^{3}$ to $10^{4}$ gates per coherence time gam17 . We can also implement spin chains experimentally using quantum dots and wells. For quantum wells built with GaAs, at room temperature the relaxation time is of the order of a few nanoseconds ohn99 . And for GaAs quantum dots, the relaxation time for a qubit is about 50 $\mu s$ at $T\approx 20mK$ han03 while for arrays of Ge/Si quantum dots we have a relaxation time of about $10\mu s$ at $T\approx 5K$ zin10 . These facts, together with the fact that in silicon quantum dots at $T\approx 1K$ we can implement single and two-qubit gates in less than $100ns$ pet22 , imply that at low temperatures ($\approx 1K$) and for silicon quantum dots it is possible to execute one hundred or so gates before the spin chain thermalizes. This number of gates is at least one order of magnitude more than what we need to implement the present proposal, where only a few gates (less than ten) is needed at a given run of the teleportation protocol.
@@
Finally, we should also contrast the present proposal with the standard way of studying QPTs from the experimental point of view. In the standard way, we must measure the one- and two-point correlation functions using, for instance, neutron scattering techniques. The non-analytic properties of these one- and two-point correlation functions at the QCP marks a QPT. In the present proposal, we need only to measure one-point correlation functions after “disturbing” the system in a specific way. Indeed, to reconstruct the state describing his qubit after the teleportation protocol (the “disturbance” in the spin chain), Bob only needs one-point correlation functions. Any single qubit density matrix is completely characterized once we know all the one-point correlation functions $\left\langle\sigma^{x}\right\rangle,\left\langle\sigma^{y}\right\rangle$, and $\left\langle\sigma^{z}\right\rangle$ nie00 . And in all the models investigated here, we only need one correlation function, $\left\langle\sigma^{z}\right\rangle$ to be specific. This is true since the qubit with Bob after the teleportation protocol is diagonal in the standard basis [see Eqs. (41)-(50)]. In other words, we are trading the measurement of two-point correlation functions to experimentally determine a QCP for the measurement of only one one-point correlation function after disturbing the system in a very particular way, namely, after applying all the gates associated with the quantum teleportation protocol.

### v2

@@ IV.2 The $\gamma$ transition
In Figs. 19 and 20 we plot, respectively, $\overline{\mathcal{F}}$ and $\overline{\mathcal{D}}$ as functions of $\gamma$ while fixing $\lambda=1.5$.
@@ V Discussion
The standard approach to study quantum phase transitions also makes use of those correlation functions or of quantities that depend on them (the magnetic susceptibility, for instance). The QCPs are detected at $T=0$ by discontinuities in the $n$-th order derivatives of those quantities at the QCPs. Quantum information theory quantities, such as bipartite or multipartite entanglement and quantum discord, are useful to spotlight QCPs too wu04 ; oli06 ; dil08 ; sar08 ; fan14 ; gir14 ; wer10 ; wer10b . See also the application of generalized phase-space techniques to detect QCPs mza19 ; mil23 . All these quantities are functions of the correlation functions of the system, the same ones we need to theoretically apply the present and the method of Ref. pav23 .
As already pointed out in Ref. pav23 , some of the standard quantities used to detect QCPs (magnetization or magnetic susceptibility, for instance) become less effective to correctly identify the QCPs with finite $T$ data wer10b . And some quantum information theory quantities, in particular the entanglement of formation woo98 , are zero at and around the QCPs as we increase $T$. These quantities cannot be employed to spotlight QCPs after a certain temperature threshold wer10b . On the other hand, a specific measure of non-classical correlations between two systems, namely, quantum discord oll01 ; hen01 , is extremely robust to spotlight QCPs at finite $T$ wer10b . The present internal teleportation based tools to detect QCPs as well as the ones of Ref. pav23 share the quantum discord robustness to detect QCPs with finite $T$ data. [^2] Note that it is important to analyze the $T=0$ limit from the finite $T$ data to make sure that no spurious classical critical point is “contaminating” the finite $T$ data. Only if the finite $T$ data consistently tend to a specific critical point as we decrease $T$ we can be reassured that we are properly seeing a QCP. See, for instance, Figs. 6, 13, 14, and 15. In addition to that, and contrary to quantum discord, the present tools and the ones of Ref. pav23 do not become computationally intractable if we deal with high spin systems (high dimensional Hilbert spaces) pav23 .
@@
Furthermore, to execute the internal teleportation protocol, and thus the present proposal, we need to be able to implement single qubit measurements, Bell state measurements, and local unitary operations on the qubits that belong to a spin chain. The single qubit measurement we need is a projective measurement onto the eigenstates of the $\sigma^{z}$ observable (our computational basis). A Bell state measurement is a joint measurement involving two qubits that aims to project the two qubits onto one of the four Bell states. This measurement can be divided into the following sequence of three steps nie00 . First, we implement a controlled-not (CNOT) gate on the pair of spins we will be measuring, now called control and target qubits. The CNOT is a two-qubit (non-local) gate whose “truth table” in the standard basis is such that it flips the spin of the target qubit if, and only if, the control qubit is given by $|1\rangle$. The standard basis can be, for instance, the eigenstates of the spin operator $\sigma^{z}$. Second, we implement a Hadamard gate on the control qubit of the CNOT gate. This is a single (local) qubit gate whose truth table is $(|0\rangle+(-)|1\rangle)/\sqrt{2}\rightarrow|0\rangle(|1\rangle)$. Third, we measure the two qubits in the standard (computational) basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Each one of the four previous outcomes in the standard basis implies that we projected the two qubits onto one of the four Bell states nie00 . The single qubit and Bell measurements are Alice’s job while Bob’s job is to measure his qubit after implementing the appropriate local operation (local gate) as given by Eqs. (8)-(11). All these gates, the ones needed by Alice and Bob, are already being implemented in spin chain-like systems using state of the art techniques ron15 ; bra19 ; xie19 ; noi22 ; xue22 ; mad22 ; xie22 . We believe that in the near future they can be implemented in the sequence described above to execute the teleportation protocol and consequently the present proposal.
@@
For instance, using superconducting qubits to realize a spin chain, it is already possible to experimentally prepare tens of spins simultaneously with coherence times around $100\mu s$. Also, single and two-qubit gates in this platform can be implemented within $10ns$ to $100ns$, leading to the execution of $10^{3}$ to $10^{4}$ gates per coherence time gam17 . We can also implement spin chains experimentally using quantum dots and wells. For quantum wells built with GaAs, at room temperature the relaxation time is of the order of a few nanoseconds ohn99 . And for GaAs quantum dots, the relaxation time for a qubit is about 50 $\mu s$ at $T\approx 20mK$ han03 while for arrays of Ge/Si quantum dots we have a relaxation time of about $10\mu s$ at $T\approx 5K$ zin10 . These facts, together with the fact that in silicon quantum dots at $T\approx 1K$ we can implement single and two-qubit gates in less than $100ns$ pet22 , imply that at low temperatures ($\approx 1K$) and for silicon quantum dots it is possible to execute one hundred or so gates before the spin chain thermalizes. This number of gates is at least one order of magnitude more than what we need to implement the present proposal, where only a few gates (less than ten) is needed at a given run of the teleportation protocol. [^3] The quantum circuit representing all the steps of the teleportation protocol can be seen in Ref. nie00 while a review on thermalization is given in Ref. mor18 . See also Ref. pek23 for a quantum circuit realization of a heat bath in solid quantum systems.
@@
Finally, we should also contrast the present proposal with the standard way of studying QPTs from the experimental point of view. In the standard way, we must measure the one- and two-point correlation functions using, for instance, neutron scattering techniques. The non-analytic properties of these one- and two-point correlation functions at the QCP marks a QPT. In the present proposal, we need only to measure one-point correlation functions after “disturbing” the system in a specific way. Indeed, to reconstruct the state describing his qubit after the teleportation protocol (the “disturbance” in the spin chain), Bob only needs one-point correlation functions. Any single qubit density matrix is completely characterized once we know all the one-point correlation functions $\left\langle\sigma^{x}\right\rangle,\left\langle\sigma^{y}\right\rangle$, and $\left\langle\sigma^{z}\right\rangle$ nie00 . And in all the models investigated here, we only need one correlation function, $\left\langle\sigma^{z}\right\rangle$ to be specific. This is true since the qubit with Bob after the teleportation protocol is diagonal in the standard basis [see Eqs. (48)-(57)]. In other words, we are trading the measurement of two-point correlation functions to experimentally determine a QCP for the measurement of only one one-point correlation function after disturbing the system in a very particular way, namely, after applying all the gates associated with the internal quantum teleportation protocol.